OpenStudy (anonymous):

The formula for degrees Celsius is: C = 9 over 5 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 4 degrees Celsius is.

OpenStudy (anonymous):

multiply by the reciprocal 5/9 to get rid of the 9/5th's and then add 32

OpenStudy (anonymous):

\[C=\frac{9}{5}(F-32)\] \[\frac{5}{9}C=\frac{5}{9}\frac{9}{5}(F-32)\] \[\frac{5}{9}C=F-32\] \[\frac{5}{9}C+32=F\]

OpenStudy (anonymous):

however i believe your formula is wrong it's suppose to be 5/9(F-32)=C

OpenStudy (anonymous):

I do believe you book or your online class is a complete fail

OpenStudy (anonymous):

lol....:)